Limx→0 (sin x -x) / tan x


1
Expert's answer
2021-06-07T03:30:22-0400

We need to calculate "lim(x\to0) \frac {(sinx-x)} {tanx}"

It is in "\frac 0 0" form after putting in place of "x" .So L Hospital's rule is applicable here .


"= lim(x\to0)\frac {\frac {d} {dx} (sinx-x)} {\frac {d} {dx}( tanx)}"


"=lim(x\to0)\frac {(cosx-1)} {(sec^2x)}"


"=\frac {(cos0 -1)} {(sec^2(0)}"


"=\frac {(1-1)} {1^2}" ( As "cos(0)=1, sec(0)=1)"


="\frac 0 1"


=0 (Ans)